An object is in simple harmonic motion with the equation of motion as x=Acos(ωt+φ),when t=T/2,the velocity of the object is......?


1
Expert's answer
2021-06-21T16:24:50-0400

Gives

"x=Acos(wt+\phi)\rightarrow(1)"

A=amplitude

Phase="\phi"

Angular frequency=w

equation (1) differencuate with respect time

Velocity

"v=Awsin(wt+\phi)"

t=0

"v=Awsin\phi\rightarrow(2)"

"t=\frac{T}{2}"

"v=Awsin(\frac{2\pi}{T}\times\frac{T}{2}+\phi)"

"v=Awsin(\pi+\phi)"

"v=Awsin\phi\rightarrow(2)"

equation (2)and(3)are same



Point P wave equation


"t=\frac{T}{2}" Velocity of object

"v=Asin\phi"